OpenStudy (anonymous):

Solve the following system. y = x + 3 2x + y = 9

OpenStudy (anonymous):

are these 2 seperate problems?

OpenStudy (anonymous):

No.

jimthompson5910 (jim_thompson5910):

2x + y = 9 2x + (x+3) = 9 ... Replace y with x+3 (since y = x+3) Now solve for x

OpenStudy (anonymous):

Hmm...

OpenStudy (anonymous):

2,5?

OpenStudy (anonymous):

I never did good on substitution or elimination.

OpenStudy (anonymous):

I just use other methods that help me. Such as this website. )

OpenStudy (anonymous):

It has for 6 months.. O.o